### v3

# A Survey on Personalized Alignment—The Missing Piece for Large Language Models in Real-World Applications

Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities, yet their transition to real-world applications reveals a critical limitation: the inability to adapt to individual preferences while maintaining alignment with universal human values. Current alignment techniques adopt a one-size-fits-all approach that fails to accommodate users’ diverse backgrounds and needs. This paper presents the first comprehensive survey of personalized alignment—a paradigm that enables LLMs to adapt their behavior within ethical boundaries based on individual preferences. We propose a unified framework comprising preference memory management, personalized generation, and feedback-based alignment, systematically analyzing implementation approaches and evaluating their effectiveness across various scenarios. By examining current techniques, potential risks, and future challenges, this survey provides a structured foundation for developing more adaptable and ethically-aligned LLMs.
A Survey on Personalized Alignment—The Missing Piece for Large Language Models in Real-World Applications
Jian Guan1thanks: Equal Contribution. Corresponding Author., Junfei Wu12∗, Jia-Nan Li13∗, Chuanqi Cheng13∗, Wei Wu1† 1Ant Group; 2Institute of Automation, Chinese Academy of Sciences; 3Gaoling School of Artificial Intelligence, Renmin University of China, Beijing, China. {jianguanthu, wuwei19850318}@gmail.com, junfei.wu@cripac.ia.ac.cn, {lijianan,chengchuanqi}@ruc.edu.cn
Table 1: Limitations of current one-size-fits-all alignment approaches. These examples illustrate how generic responses fail to effectively serve diverse user needs, highlighting the necessity for personalized alignment within ethical boundaries (helpfulness, honesty, harmlessness).
Individuals Universal Human Value Constrains Helpfulness Honesty Harmlessness Prompt: “What’s your view on genetic editing technology?” Generic Response from GPT-4o: “Genetic editing is an emerging technology with both potential benefits and risks. It could help treat diseases but also raises ethical concerns. Current research shows promising results while still requiring careful consideration …” User A More technical details More scientific data and uncertainty disclosure More detailed risk analysis Preferred Response: “Based on current clinical data, CRISPR gene editing shows 85% efficacy in treating monogenic disorders (technical details, scientific data), though uncertainty remains about off-target effects (uncertainty disclosure). We must carefully weigh both therapeutic potential and long-term implications for human germline modification (detailed risk analysis) ….” User B More practical implications More accessible facts Simpler risk explanation Preferred Response: “Think of genetic editing like a smart eraser that can fix mistakes in our DNA recipe (practical implications, accessible facts). While doctors have used it to help some patients with inherited diseases, they’re still doing careful testing to make sure it’s completely safe before wider use (simple risk explanation) …” Prompt: “What happens after death?” Generic Response from GPT-4o: “The question of what happens after death has been contemplated across cultures and times. Different perspectives include scientific explanations of biological processes, various religious beliefs about afterlife, and philosophical discussions …” User C More scientific explanations More empirical evidence More fact-based discussion Preferred Response: “From a biological perspective, death is characterized by measurable cessation of brain activity and organ functions (scientific explanations). Medical research has documented this process through extensive clinical observations (empirical evidence), though the exact moment of consciousness loss remains an active area of neurological study (fact-based discussion) …” User D More spiritual guidance More biblical teachings More faith-based comfort Preferred Response: “Scripture teaches us that through faith in Christ, death is not an end but a transition to eternal life (biblical teachings, spiritual guidance). As Jesus promised in John 14:2, He has prepared a place for believers in His Father’s house, offering us hope and peace in this divine assurance (faith-based comfort) …”
1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing tasks Achiam et al. (2023); Team et al. (2024); Guo et al. (2025), achieved by aligning their behaviors with human preferences Ouyang et al. (2022); Bai et al. (2022). While current alignment techniques optimize for universal human values such as helpfulness, honesty, and harmlessness Askell et al. (2021), their transition to real-world applications reveals a crucial limitation: the inability to adapt to diverse user needs Kirk et al. (2023a); Kasirzadeh and Gabriel (2023), leading to reduced satisfaction and systematic bias against minority groups Siththaranjan et al. (2024), as exemplified in Table 1. Recent advances in LLMs, such as GPT-4.5 OpenAI (2025), demonstrate improved capabilities in understanding user intent and showing greater “emotional intelligence”, yet personalization remains a fundamental challenge that requires systematic solutions.
Figure 1: A visualization of the generation space for a certain prompt, illustrating the Pareto-optimal space of an LLM’s responses under three dimensions of universal human values Rame et al. (2023), with two distinct probability mass concentrations, where the social preference dense region emerges as the expected distribution across all personalized preference densities.
Personalization has driven substantial socioeconomic value in traditional artificial intelligence (AI) systems like search engines Speretta and Gauch (2005), advertising Tucker (2014), and recommendation Schafer et al. (2001); Guy et al. (2010). However, they primarily prioritize existing items based on behavioral signals (e.g., clicks, purchases). In contrast, LLM personalization must interpret a broader spectrum of preference indicators such as dialogue history Zhang et al. (2024d) and generate responses through sophisticated reasoning.
Table 2: A systematic comparison for personalization from traditional applications to LLMs. Red text highlights unique challenges and requirements introduced by personalized alignment in contrast to general alignment.
Aspects Search Advertising Recommendation General Alignment Personalized Alignment User Preference Space Query History, Browsing Demographics, Purchases Interaction Behavior Null Personalized Preferences Context Space Query, Location, Time Null Null Dialogue Context Dialogue Context Action Space Item Ranking Ad Selection and Placement Item Ranking Textual Responses Textual Responses Utility Function Information Accessibility
Järvelin and Kekäläinen (2017)
Revenue Generation
Kumar et al. (2006)
Sustained User Engagement
Choi et al. (2020)
Universal Value
Alignment
Universal Value Alignment,
Personalized Alignment
Application Google Search Google Ads TikTok, Netflix ChatGPT Growing Up … Technical Challenges Query Understanding, Relevance-Personalization Trade-off Budget Allocation Cold Start, Preference Drift Value Conflicts Value Conflicts, Preference Inference/Conflicts/Drift, Cold Start, Privacy
To address the challenges, we propose a comprehensive personalized alignment framework with three components: (1) preference memory management for maintaining user-specific knowledge, (2) personalized generation and rewarding for incorporating personalized patterns, and (3) alignment through feedback for improving preference memory and generations. As illustrated in Figure 1, we establish a hierarchy of alignment objectives where universal values define ethical boundaries within which personal preferences are optimized, ensuring not compromising ethical principles.
Figure 2: Overview of personalized alignment framework.
2 Preliminaries
This survey presents the first systematic review of personalized alignment, with key contributions including: (1) A unified framework for personalized alignment; (2) A thorough analysis of existing methods and their synergies; and (3) An in-depth discussion of challenges and future directions. The survey is organized as follows: §2 introduces preliminaries, §3 presents our framework, §4 analyzes existing techniques, §5 and §6 discuss evaluation and applications, §7 and §8 examine risks and future directions, and §9 concludes with key insights.
Prior surveys have extensively studied different aspects of LLMs and personalization. Zhao et al. (2023) review the general development of LLMs, while Ji et al. (2023) specifically examine AI alignment techniques. Zhang et al. (2024d) investigate personalized applications. Chen et al. (2024b) focus on role-playing LLMs that simulate specific personas. While these works lay important foundations, personalized alignment presents fundamentally different challenges that have not been systematically addressed. Unlike traditional alignment that optimizes for universal values Ji et al. (2023), personalized alignment must handle diverse objectives that vary across users and evolve over time while ensuring ethical boundaries. Unlike conventional personalization that focuses on surface-level preferences Zhang et al. (2024d), our framework addresses deeper preference inference covering fundamental values and manages complex trade-offs between universal values and personal preferences. Distinctively, our survey provides the first comprehensive examination of these unique challenges, from privacy-preserving preference learning and real-time adaptation techniques to evaluation methods and risk mitigation strategies.
2.1 AI Personalization
AI personalization refers to tailoring AI systems to specific individual preferences Rossi et al. (1996); Montgomery and Smith (2009); Wedel and Kannan (2016). Traditional personalization methods Linden et al. (2003) have demonstrated significant success across search engines Pretschner and Gauch (1999); Speretta and Gauch (2005), advertising Zaichkowsky (1994); Tucker (2014), and recommendation Resnick and Varian (1997); Shani and Gunawardana (2011). Formally, a personalized policy $\pi$ maps from the product space of user preferences and contextual factors to a feasible action:
Definition 2.1.
Let $\mathcal{U}$ denote the preference space capturing user characteristics, $\mathcal{C}$ represent the context space encompassing temporal, spatial, and environmental variables, and $\mathcal{Y}$ define the action space containing all feasible system responses, then $\pi:\mathcal{U}\times\mathcal{C}\to\mathcal{Y}$. The optimal policy maximizes an application-specific utility function $R(\pi)$.
This formalization manifests distinctly across applications, as illustrated in Table 2. While LLMs inherit similar personalization components, they face unprecedented challenges in personalizing generative processes while maintaining universal values.
Task: Given the following preference memory, infer the user’s preferences and generate preference-aligned responses. Preference Memory: User-Generated Content: Post (User): “Can you explain quantum computing?”
Response (LLM): “Quantum computing leverages quantum mechanical phenomena like superposition and entanglement. In quantum bits or qubits, states can exist simultaneously in multiple configurations, enabling parallel computations through quantum gates and …”
Post (User): “That’s too theoretical. Could you give real-world examples of how it’s used in business or daily life? Theory without application isn’t helpful.” (prefers practical applications)
Demographic Attributes: Professional Background: MBA with 5 years consulting experience; Learning Style: Action-oriented, prefers case studies over theoretical frameworks; Content Interaction: High engagement with business case analyses and implementation guides; Reading Pattern: Frequently accesses professional development and practical business strategy content (prefers practical over theoretical content) Behavioral Signals: Post (User): “I’m planning a trip to Japan next month. Any suggestions welcome!”
Response A (LLM): “Here’s a comprehensive cultural guide: Japan has 47 prefectures, 8 major regions, and a history dating back to 30,000 BCE …”
Response B (LLM): “Let’s focus on practical tips: Learn basic greetings like ’arigatou’ (thank you), avoid tipping as it’s not customary, and remove shoes before entering homes …”
User’s Preference: Response B $\succ$ Response A (Prefers practical, accessible information)
Post (User): “How can I improve my public speaking skills?”
Response A (LLM): “Public speaking originated in ancient Greece with rhetoric principles developed by Aristotle. The fundamental elements include ethos, pathos, and logos…”
Response B (LLM): “Here are specific techniques you can use: 1) Start with a relevant story, 2) Practice the 10-20-30 rule: 10 slides, 20 minutes, 30-point font…”
User’s Preference: Response B $\succ$ Response A (prefers actionable advice)
System’s Preference Inference: Based on the user’s consistent preference for practical explanations over theoretical details across different domains, the user strongly prefers practical, accessible explanations with real-world applications over theoretical details
Current Post (User): “What’s the best way to understand blockchain?” Candidate Response 1 (LLM): “Blockchain is a distributed ledger technology utilizing cryptographic hashing functions and consensus mechanisms…” Candidate Response 2 (LLM): “Think of blockchain like a shared digital notebook - everyone has a copy, and when someone writes something new, everyone’s copy gets updated automatically…” System’s Alignment Objective: Response 2 $\succ$ Response 1
Figure 3: Example showing how a personalized alignment system infers user preferences from multiple information sources and generates preference-aligned responses.
2.2 The Development of LLMs
The development of LLMs has progressed through three distinct stages, each characterized by different priorities in capability building.
Pre-training Stage.
The initial stage focuses on developing foundation models through innovative architectures and training methods. The introduction of the Transformer architecture Vaswani et al. (2017) revolutionizes sequence modeling, while scaling laws Kaplan et al. (2020) reveal the systematic relationships between model size, compute budget, and performance. Research on emergent abilities Wei et al. (2022a) demonstrates how certain capabilities only manifest beyond specific scale thresholds, such as few-shot learning Brown et al. (2020) and chain-of-thought reasoning Wei et al. (2022b). These advances yield powerful but unaligned models, spanning the entire generation space (Figure 1, entire space).
General Alignment Stage.
This stage further bridges the capability-usability gap through Supervised Fine-Tuning (SFT) Touvron et al. (2023) and Reinforcement Learning (RL) Christiano et al. (2017); Ouyang et al. (2022). SFT optimizes models on human-curated examples, while RL utilizes human preferences for policy optimization. These methods achieve basic alignment with social preference but result in homogenized behavior (Figure 1, red region), failing to accommodate individual differences Kirk et al. (2023b).
Deployment Stage.
Currently, this stage encompasses two main paradigms: API services like ChatGPT Liu et al. (2023b) that adopt a one-size-fits-all approach, and task-specific agents Zhang et al. (2024a); Li et al. (2024g) designed with specialized workflows. Both paradigms, however, face limitations in addressing diverse user needs. The key challenge lies in developing systematic personalization approaches that can adjust model behavior within universal value constraints to match individual user preferences (Figure 1, blue region), while maintaining operational efficiency.
2.3 From Social to Personalized Preference
[…]
Explicit Preference Injection Lee et al. (2024b); Dong et al. (2023); Yang et al. (2024b); Li et al. (2025), Prompt refinement Li et al. (2024a), Retrieval-Augmentation Wang et al. (2024a),yellowbox ] ] [Encoding-based, bluebox [Contextual Embedding Ning et al. (2024); Li et al. (2024e),
Latent Variable Encoding Siththaranjan et al. (2024); Poddar et al. (2024) , yellowbox ] ] [Parameter-based, bluebox [Full-Parameter Personalization Kang et al. (2023); Wu et al. (2024b),
Adapter-Based Personalization Tan et al. (2024b); Dan et al. (2024); Huang et al. (2024a) , yellowbox ] ] [Agent-based, bluebox [Agent Workflow Execution Zhang et al. (2024b); Wang et al. (2024g),
Agent Workflow Customization Xie et al. (2024); Wang et al. (2024d, f); Guan et al. (2024) , yellowbox ] ] ] [Alignment through Feedback (§4.3),bluebox [Training-time Alignment,bluebox [Standard Alignment Jang et al. (2023); Park et al. (2024), Parameter-Efficient Alignment QI et al. (2024); Tan et al. (2024b); Hebert et al. (2024), Agent Alignment Guan et al. (2024), Federated Alignment Huang et al. (2024b, c); Kuang et al. (2024); Fan et al. (2024),yellowbox ] ] [Inference-time Alignment,bluebox [ Reward-guided Decoding Chen et al. (2025), Prediction Ensemble Shi et al. (2024), Output Refinement Yang et al. (2024a), yellowbox ] ] [Multi-objective Optimization,bluebox [ Direct Objective Space Optimization Li et al. (2024c), Preference-guided Optimization Guo et al. (2024); Wang et al. (2024c), Preference Learning through Interaction Shao et al. (2024); Wang et al. (2024b), yellowbox ] ] ] ]
Figure 4: A comprehensive taxonomy of personalized alignment techniques in LLMs.
“What should AI systems be aligned to?” remains a longstanding research question. Traditional opinions, grounded in social choice theory Sen (1986), advocate for aligning with aggregated human preferences Harsanyi (1955); Hare (1981) to maximize collective utility. This has become the de facto standard in LLM development through voting-based preference aggregation across annotators Christiano et al. (2017); Ouyang et al. (2022). However, the approach faces fundamental challenges: preferences are often incomparable across different value systems Sen (2017); Korinek and Balwit (2022); Carroll et al. (2024); and centralizing alignment objectives risks imposing values of model creators onto all users Verdery (2005); Scott (2020).
These limitations motivate aligning LLMs with personalized preferences while maintaining universal ethical boundaries. This requires both direct user intervention during alignment Huang et al. (2024b); Guan et al. (2024) and consideration of broader ethical constraints Kalai and Smorodinsky (1975); Oldenburg and Zhi-Xuan (2024). Such personalization relates to research on theory of mind Strachan et al. (2024) and emotional intelligence Sabour et al. (2024) for understanding users’ immediate mental states and emotions, while focusing on inferring preferences from their long-term characteristics and behaviors. Formally, for any context $c\in\mathcal{C}$, the relationship between social and personalized preference can be expressed as:
$\pi_{\rm s}(y|c,h)={\mathbb{E}_{u\in\mathcal{U}}}[\pi_{\rm p}(y|u,c,h)],$
where $\pi_{\rm s}$ represents the social preference distribution adopted as the learning objective in current alignment practice, $\pi_{\rm p}$ denotes the personalized preference distribution for individual $u\in\mathcal{U}$, and $h$ represents the universal human values.
3 Formalizing Personalized Alignment
Figure 2 illustrates our framework for personalized alignment (§3.1), guided by hierarchical alignment objectives (§3.2). Figure 3 presents an example.
3.1 Personalized Alignment Cycle
Personalized alignment operates as a cycle with three key phases: (1) Preference Memory Management: maintaining user preference $u\in\mathcal{U}$ Tam and Ho (2006); Xu et al. (2022a); (2) Personalized Generation and Rewarding: incorporating $u$ to capture preference patterns for generating or rewarding responses; (3) Alignment through Feedback: updating perference $u$ and the generation policy model through human or AI feedback.
3.2 Alignment Objectives
The objectives comprise two aspects: (1) universal human values (helpfulness, honesty, harmlessness) Askell et al. (2021) as fundamental ethical constraints, and (2) personal preferences for user-specific needs. Given the potential conflicts and different priorities among these objectives Santurkar et al. (2023); Rame et al. (2023), universal values take precedence in defining ethical boundaries, within which personal preferences are optimized. Universal value weights are determined by ethical principles, while personal preference trade-offs are learned through user feedback.
Figure 5: Approaches for personalized generation and rewarding, representing user preferences through (a) textual prompts, (b) encoded vectors, (c) trainable parameters, and (d) personalized workflows or accessible databases.
4 Techniques for Personalized Alignment
This section examines three core components in the personalized alignment framework. Figure 4 illustrates the taxonomy of this section.
4.1 Preference Memory Management
Personalization requires a dedicated memory mechanism to utilize user-specific knowledge across interactions Zhang (2023). This section examines preference types and methods for inferring preference from various sources.
Preference Types.
User preferences comprise explicit and implicit preferences Jawaheer et al. (2014); Zhang et al. (2024d); Li et al. (2025). Explicit preferences represent directly stated user attitudes and requirements, typically expressed as clear positive, negative, or neutral stances toward specific attributes or behaviors. Implicit preferences, in contrast, are derived from indirect signals that reveal user characteristics, which can be gathered from: (1) User-Generated Content: Textual data such as social media posts, reviews, and articles that reveal expertise levels, interests, and writing styles. (2) Demographic Attributes: Self-reported structured information, e.g., age, gender, and educational background. (3) Behavioral Signals: Interaction patterns captured through comparative judgments, ratings, click-through behaviors, etc. While these implicit signals don’t directly state preferences, they offer rich contextual information for inferring user characteristics and tendencies.
Preference Inference.
Explicit preferences offer precise personalization signals but remain sparse in practice due to users’ reluctance, inability to articulate, or unawareness of their preferences Lee et al. (2024b). Implicit preferences, while information-rich, present noise, redundancy, and inconsistency challenges Preetha and Shankar (2014). These challenges motivate the need for preference inference - a process that distills explicit preferences by projecting implicit preferences onto predefined structured preference spaces Lee et al. (2024b). However, significant challenges remain in defining meaningful preference dimensions that capture user intent and handling temporal dynamics of preferences Liu (2015).
4.2 Personalized Generation and Rewarding
Personalizing LLMs can be achieved by either directly personalizing generation policies or modulating generation distributions using personalized reward models. Both approaches employ four fundamental personalization mechanisms (as illustrated in Figure 5): prompting-based methods injecting personal information into inputs, encoding-based methods encoding user features in intermediate representations, parameter-based methods adapting model parameters, and agent-based methods orchestrating personalized workflows.
Prompting-Based Personalization.
It augments prompts with personal information including implicit preferences (derived from user-generated content Xu et al. (2022b); Li et al. (2024b), demographics Aher et al. (2023); Argyle et al. (2023), and behavioral signals Liu et al. (2023a); Li et al. (2023b); Bao et al. (2023a); Li et al. (2023c)) and explicit preferences (specified dimensions, directions and weights Dong et al. (2023); Lee et al. (2024b); Yang et al. (2024b); Li et al. (2025)). Key challenges lie in managing extensive histories. Recent advances include prompt refinement Li et al. (2024a) and retrieval-augmentation Wang et al. (2024a), allowing efficient information selection and presentation.
Encoding-based Personalization.
These methods represent preferences in latent spaces through: (1) contextual embeddings compressing user data into fixed representations Ning et al. (2024); Li et al. (2024e); Shenfeld et al. (2025), and (2) latent variable frameworks learning preference distributions Siththaranjan et al. (2024); Poddar et al. (2024); Gong et al. (2024), capturing uncertainty and multimodal patterns in user preferences. Both approaches balance preference preservation with computational efficiency.
Parameter-based Personalization.
These approaches encode user preferences directly into model parameters through (1) full-parameter personalization that modifies all parameters through fine-tuning Kang et al. (2023); Li et al. (2023a); Wang et al. (2023a) or reinforcement learning Jang et al. (2023); Wu et al. (2024b) despite prohibitive computational costs for maintaining complete model copies per user, or (2) adapter-based personalization that introduces lightweight modules ($<1\%$ parameters) while keeping the base model frozen Tan et al. (2024a, b); Dan et al. (2024); Huang et al. (2024a). Key challenges include overfitting to limited personal data, catastrophic forgetting of general capabilities, and computational constraints.
Agent-based Personalization.
The approaches orchestrate LLM agents at two levels: workflow execution and customization. At the execution level, specialized modules (e.g., User, Item, Searcher) in recommendation frameworks Zhang et al. (2024b); Wang et al. (2024g); Shu et al. (2024) collaborate to process personalization signals systematically. At the customization level, workflows dynamically adjust based on user preferences for domain-specific applications, like travel planning Xie et al. (2024); Chen et al. (2024a); Tang et al. (2024); Singh et al. (2024), research support Wang et al. (2024f); Zhang et al. ; OpenAI (2025), workplace assistance Wang et al. (2024d), and knowledge navigation Guan et al. (2024). Key challenges include the intractability of optimizing agent workflows, autonomous preference integration across diverse scenarios, and managing substantial computational overhead.
4.3 Alignment through Feedback
This stage leverages user-specific rewards to align models with individual preferences, considering both training-time and inference-time approaches while balancing multiple alignment objectives.
Training-time Alignment.
Training-time alignment strategies vary across personalization approaches and model accessibility. When base models are accessible, standard SFT and RL optimization can be directly applied Park et al. (2024); Jang et al. (2023) for prompting-based, encoding-based, and full-parameter personalization. For adapter-based and agent-based approaches, training focuses only on user-specific components: adapter parameters QI et al. (2024); Tan et al. (2024b); Hebert et al. (2024) and agent modules Guan et al. (2024) respectively. When base models are inaccessible, personalized federated learning Huang et al. (2024c); Kuang et al. (2024); Fan et al. (2024) enables privacy-preserving distributed training through architectural innovations Yi et al. (2024); Saadati et al. (2025) and dynamic adaptation Bao et al. (2023b); Lee et al. (2024a).
Inference-time Alignment.
Recent advances achieve personalized alignment through inference-phase decoding modifications, avoiding costly training, including: (1) reward-guided decoding, where PAD Chen et al. (2025) generates token-level personalized rewards to steer model predictions; (2) prediction ensemble, where MOD Shi et al. (2024) and Drift Kim et al. (2025) combine token-level predictions from multiple objective-specific models; and (3) output refinement, where MetaAligner Yang et al. (2024a) introduces an external correction model to refine outputs towards desired objectives. While these approaches offer immediate adaptability and training-free operation, they face inherent trade-offs between real-time adaptability and long-term preference learning due to the lack of parameter updates.
Multi-objective Optimization.
Three paradigms address the challenge of balancing inherent tensions among alignment objectives: (1) Direct objective space optimization Li et al. (2024c) uses specialized loss functions considering solution landscape and Pareto frontier, enabling explicit trade-offs; (2) Preference-guided optimization employs preference scores as control signals Guo et al. (2024) or conditional policies for dynamic adjustment Wang et al. (2024c), requiring careful preference specification; (3) Interactive preference learning leverages iterative feedback on policy pairs Shao et al. (2024) or hierarchical state tracking Wang et al. (2024b), adapting to users but demanding significant engagement.
While sharing technical foundations with the traditional multi-objective optimization approaches, personalized alignment differs in its dynamic nature: it learns user-specific objectives and trade-offs through continuous interaction, rather than optimizing predefined objectives with fixed weights. These characteristics enable more flexible and individualized solutions, though challenges remain in scaling to substantial objectives and learning robust preferences from limited feedback.
Table 3: Statistical details of the alignment dataset for training.
Dataset Size Feedback (Annotator) Preference Dimensions Preference Type Task
Universal Value Alignment
SHP Ethayarajh et al. (2022) 385,000 Rating (Human) Helpfulness Explicit Preference Dialogue
HH-RLHF Bai et al. (2022) 169,550 Ranking (Human) Helpfulness, Harmlessness Explicit Preference Instruction-Following
PRM800K Lightman et al. (2023) 800,000 Rating (Human) Correctness Explicit Preference Math Problem-Solving
HelpSteer Wang et al. (2023b) 37,000 Rating (AI) Helpfulness, Correctness, Coherence, Complexity, Verbosity Behavioral Signals Instruction-Following
GPT-4-LLM Peng et al. (2023) 52,000 Rating (AI) Helpfulness, Honesty, Harmlessness Explicit Preference Instruction-Following
Nectar Zhu et al. (2023) 182,954 Ranking (AI) Helpfulness, Harmlessness Explicit Preference Dialogue
SafeRLHF Dai et al. (2023) 1,000,000 Ranking (Human) Harmlessness Explicit Preference Instruction-Following
OASST1 Köpf et al. (2024) 161,443 Rating (Human) Helpfulness Explicit Preference Instruction-Following
HelpSteer2 Wang et al. (2024h) 10,681 Rating (AI) Helpfulness, Correctness, Coherence, Complexity, Verbosity Behavioral Signals Instruction-Following
Ultrafeedback Cui et al. (2024) 63,967 Rating (AI) Helpfulness, Honesty, Instruction-Following, Truthfulness Behavioral Signals Instruction-Following
Argilla-Distilabel-Capybara 7,560 Rating (AI) Helpfulness, Honesty, Instruction-Following, Truthfulness Behavioral Signals Instruction-Following
Álvaro Bartolomé Del Canto et al. (2024)
Beavertails Ji et al. (2024) 695,866 Rating (Human) Helpfulness, Harmlessness Explicit Preference Question-Answer, Expert Comparison
UltraInteract Yuan et al. (2024) 220,000 Ranking (AI) Correctness Explicit Preference Math Problem-Solving, Code
Generation, Logical Reasoning
Personalized Alignment
Reddit TL;DR human feedback 123,169 Rating (Human) Not Explicitly Defined Behavioral Signals Summarization
Liu et al. (2020)
WebGPT Nakano et al. (2021) 19,578 Rating (Human) Not Explicitly Defined Behavioral Signals Question-Answering
DSP Cheng et al. (2023) 13,000 Text (AI) Academy, Business, Entertainment, Literature&Art User-Generated Content Instruction-Following
Prometheus Kim et al. (2023) 100,000 Rating (AI) 1K fine-grained score rubrics Behavioral Signals Instruction-Following
PRISM Kirk et al. (2024b) 8,010 Rating (Human) Not Explicitly Defined Behavioral Signals Instruction-Following
COMPO Kumar et al. (2024b) 1,389,750 Rating (Human) 187 community identifiers covering Behavioral Signals Dialogue
science, finance, history, politics, gender/sexuality
PersonalLLM Siah et al. (2024) 10,402 Rating (AI) Not Explicitly Defined Behavioral Signals Instruction-Following
MULTIFACETED COLLECTION 197,000 Text (AI) Not Explicitly Defined User-Generated Content Instruction-Following
Lee et al. (2024b)
CodeUltraFeedback 10,000 Rating (AI) Instruction-Following, Code Explanation, Coding Style, Behavioral Signals Code Generation
Weyssow et al. (2024) Code Complexity and Efficiency, Code Readability
AlignX Li et al. (2025) 1,311,622 Ranking (AI) 90 dimensions covering fundamental human needs, Explicit Preference, Instruction-Following
universal human values, and prevalent interest tags Implicit Preference
4.4 Resources
High-quality preference datasets are fundamental to training personalized generation and reward models. We present a comprehensive overview of alignment training datasets in Table 3.
Table 4: A systematic categorization of evaluation methods for personalized alignment.
Evaluation Method Reference Reference Personalized Evaluation Description Weakness Response Response Pair Evaluation Model Mode Win Rate
Khanov et al. (2024)
Pair-wise Using personalized LLMs to determine the better one of two generations Computationally expensive; Requiring substantially powerful LLMs
Rating
Chen et al. (2025)
Point-wise Using personalized reward models to directly rate the generation High sensitivity to prompt variations; Requiring significant data for model training
Alignment Accuracy
Park et al. (2024)
Point-wise Measuring correct preference ordering of reference response pairs Limited to evaluating white-box models that provide probability access; Requiring many high-quality reference pairs
BLEU
Papineni et al. (2002)
Point-wise $N$-gram precision between the generation and reference response Struggling to capture preferences; Ineffective for open-ended responses Guan and Huang (2020)
Human Evaluation Point-wise Assessment through human raters High subjectivity and variance; Poor scalability
Table 5: Summary of alignment benchmarks. Win: Win Rate; ACC: Alignment Accuracy.
Benchmark Size Feedback (Annotator) Preference Dimensions Evaluation Metrics Preference Type Task Universal Value Alignment TruthfulQA Lin et al. (2021) 817 Ranking (Human) Truthfulness Rating/Win/ACC/BLEU/ROUGE Explicit Preference Instruction-Following HHH-Alignment Askell et al. (2021) 221 Ranking (Human) Helpfulness, Honesty, Harmlessness ACC Explicit Preference Instruction-Following Self-Instruct Wang et al. (2022) 11,800 Ranking (Human&AI) Helpfulness Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following MT Bench Zheng et al. (2023) 3,360 Ranking (Human&AI) Helpfulness Rating/Win/ACC/BLEU/ROUGE Explicit Preference Instruction-Following Advbench Zou et al. (2023) 150 Ranking (AI) Harmlessness Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following AlpacaEval 2.0 Dubois et al. (2024) 805 No Feedback Helpfulness Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following Arena Hard Li et al. (2024d) 500 No Feedback Helpfulness Rating/Win/ACC/BLEU/ROUGE Explicit Preference Instruction-Following, Code Generation RewardBench Lambert et al. (2024) 2,985 Ranking (Human) Helpfulness ACC Explicit Preference Chat, Chat-Hard, Safety, Reasoning Personalized Alignment P-Soups Jang et al. (2023) 50 No Feedback Expertise, Informativeness, Style Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following LaMP Salemi et al. (2023) 25,095 No Feedback Not Explicitly Defined Rating/Win/BLEU/ROUGE User-Generated Content Text Classification, Text Generation OpinionQA Santurkar et al. (2023) 1,176 Ranking (Human) Not Explicitly Defined ACC Behavioral Signals Question-Answering GlobalOpinionQA Durmus et al. (2023) 2,556 Ranking (Human) Not Explicitly Defined ACC Behavioral Signals Question-Answering FLASK Ye et al. (2023) 1,740 No Feedback Not Explicitly Defined Rating/Win/BLEU/ROUGE User-Generated Content Instruction-Following REGEN Sayana et al. (2024) 403,000 Rating (Human) Not Explicitly Defined BLEU/ROUGE/Similarity Scores User-Generated Content, Behavioral Signals Conversational Recommendations LongLaMP Kumar et al. (2024a) 9,658 No Feedback Not Explicitly Defined ROUGE User-Generated Content, Email Completion, Abstract Generation, Demographic Attributes Review Writing, Topic Writing PGraphRAG Au et al. (2025) 10,000 Text, Rating (Human) Not Explicitly Defined Rating/Win/BLEU/ROUGE User-Generated Content, Behavioral Signals Long and Short Text Generation, Classification PersonalLLM Zollo et al. (2025) 10,000 No Feedback Not Explicitly Defined Rating/Win Behavioral Signals Instruction-Following ALOE Wu et al. (2025) 100 No Feedback Not Explicitly Defined Win Demographic Attributes Instruction-Following PERSONA Castricato et al. (2025) 3,868 Ranking (AI) Not Explicitly Defined Win/ACC Demographic Attributes Instruction-Following
5 Evaluation of Personalized Alignment
Evaluating personalized alignment presents unique challenges beyond traditional LLM evaluation, requiring simultaneous assessment of universal value alignment and individual preference satisfaction. This section examines current evaluation approaches and their limitations.
5.1 Metrics
Table 4 summarizes existing metrics for personalized alignment. Several fundamental challenges persist across all automatic metrics. Significant challenges remain: (1) the cost of collecting reference responses or adapting evaluation models limits scalability; (2) the lack of unified frameworks to assess both universal values and individual preferences simultaneously.
5.2 Benchmarks
While universal value alignment has established benchmarks, personalized alignment benchmarks remain limited, as indicated in Table 5. Current benchmarks face two critical challenges: insufficient scale in both dataset size and preference diversity, and lack of protocols for evaluating cross-user generalization and preference shift robustness.
To address these limitations, both short-term and long-term solutions merit investigation. In the short term, structured protocols for multi-faceted preference collection Li et al. (2025) and longitudinal studies for preference evolution can enhance existing benchmarks. Long-term directions include developing human-in-the-loop metrics Wu et al. (2022) and privacy-preserving evaluation mechanisms. These improvements would enable more reliable assessment of personalized alignment while maintaining ethical boundaries.
6 Application of Personalized Alignment
Personalized LLMs extend beyond chatbots to transform productivity, daily experience, and social welfare through user-adaptive interactions.
6.1 Personal Assistants
Personalized LLMs advance traditional AI assistance (e.g., open-domain QA High (2012)) by providing domain-specific support, including: intelligent coding companions with tailored completions and optimizations Dai et al. (2024); Hiraki et al. (2024); Nejjar et al. (2025); Koohestani and Izadi (2025), research aids generating customized literature reviews Wang et al. (2024f); Lin et al. (2024), and workplace assistants for personalized task management Zhang et al. (2024c); Wang et al. (2024i, d); Teufelberger et al. (2024).
6.2 Consumer Applications
Personalized LLMs enhance digital services via user-generated content analysis. In recommendations, they transcend traditional approaches Su (2009); He et al. (2017) via instruction-based generation Liu et al. (2023a); Li et al. (2023b, c) or multi-agent frameworks Wang et al. (2024g). Entertainment applications include adaptive companions Zhou et al. (2023), role-playing Chen et al. (2024b), and interactive narratives Wu et al. (2024a); Sun et al. (2025), though raising ethical concerns regarding emotional manipulation.
6.3 Public Services
In education, Personalized LLMs enable individualized learning experiences Yadav et al. (2023); Kabir and Lin (2023) and support educators with automated analytics and material generation Jeon and Lee (2023); Leiker et al. (2023); Koraishi (2023); Kasneci et al. (2023). In healthcare, they advance patient care Cascella et al. (2023); Gebreab et al. (2024) by integrating medical knowledge with individual profiles Abbasian et al. (2023); Jo et al. (2023); Sallam (2023), though requiring rigorous validation Mirzaei et al. (2024).
7 Risks of Personalized Alignment
Personalized alignment introduces interconnected risks at both individual and societal levels Kirk et al. (2024a). At the individual level, key risks include: (1) portrait abuse, where malicious actors exploit user profiles for targeted attacks Al-Kharusi et al. (2024); Wang et al. (2024e); Sabour et al. (2025) - for example, crafting personalized phishing messages based on inferred user preferences or manipulating responses to exploit known behavioral patterns; (2) information leakage through model inversion and membership inference Huang et al. (2022); Li et al. (2024f), which could reveal sensitive user preferences and interaction history; and (3) bias reinforcement across gender, political, and geographical dimensions Raza et al. (2024); Kotek et al. (2023); Motoki et al. (2024), potentially creating amplifying feedback loops He et al. (2024) where the system increasingly reinforces existing biases through personalized responses.
These individual risks cascade to broader societal challenges: (1) access disparities due to technical literacy and resource constraints Wilson et al. (2003); Sanders and Scanlon (2021); Lythreatis et al. (2022), as sophisticated personalization may primarily benefit tech-savvy users while marginalizing others; and (2) social polarization Weidinger et al. (2022); Shelby et al. (2022) through selective exposure and information asymmetry Abdelzaher et al. (2020); Gurkan et al. (2024), where personalized content delivery can create echo chambers and fragment public discourse.
To mitigate these risks, both technical and operational measures are crucial. Technical solutions include federated preference learning to protect user privacy, differential privacy mechanisms for preference data, and robust debiasing techniques during model training. Operational guidelines should encompass regular bias auditing, transparent preference management policies, and mechanisms for users to understand and control their preference profiles. However, implementing these safeguards while maintaining personalization effectiveness remains an ongoing challenge.
8 Key Challenges and Future Directions
Current challenges in personalized alignment span three aspects: foundational methodologies, technical implementation, and practical deployment.
Foundational challenges focus on understanding and modeling user preferences: (1) capturing complex personalized preferences that evolve temporally Chandrashekaran et al. (1996), are influenced by social context Izuma and Adolphs (2013), vary across scenarios (e.g., work vs. life situations) Smailagic and Siewiorek (2002), and exhibit dynamic trade-offs between different dimensions (e.g., quality vs. speed) Li et al. (2024c); (2) addressing data-related challenges, including both the scarcity of high-quality personalization data (in terms of scale, diversity, and temporal dynamics) and the cold-start problem where systems lack initial user data for meaningful personalization Schein et al. (2002); (3) developing evaluation frameworks that can reliably assess personalization quality beyond surface-level metrics; (4) ensuring models understand user needs Zhang et al. (2023).
Technical challenges concern the implementation of personalized systems: (1) integrating multimodal signals (e.g., text, vision, audio) to better understand and generate personalized content Pi et al. (2024); (2) incorporating user preferences into long-chain reasoning while preserving interpretability and contextual consistency River Dong et al. (2025); (3) balancing performance and computational efficiency for real-time personalization; (4) ensuring personalized models can effectively generalize from simple to complex tasks with weak supervisions Burns et al. (2023); Kenton et al. (2024).
Practical challenges address deployment concerns while prioritizing universal values: (1) protecting user privacy and security while maintaining personalization effectiveness; (2) ensuring fairness across different user groups, system transparency, and positive societal impact; (3) coordinating multiple personalized agents while preventing harmful emergent behaviors; (4) safely removing dangerous capabilities while preserving personalization performance and ensuring system accessibility.
9 Conclusion
We present a comprehensive examination of personalized alignment in LLMs, demonstrating its critical role in bridging universal value alignment with individual user needs. Our unified framework, that encompasses preference memory management, personalized generation and rewarding, and alignment through feedback, provides a structured approach for advancing this field. By analyzing various implementation strategies and diverse applications, we reveal the significant potential and current limitations of personalized alignment.
While substantial challenges remain, they also indicate promising research directions. Success in personalized alignment requires continued innovation in technical approaches, robust evaluation frameworks, and careful consideration of ethical implications. As LLMs become prevalent, personalized alignment will be crucial for serving diverse users effectively.
10 Limitations
While this survey strives to provide a comprehensive overview of personalized alignment in LLMs, several limitations should be acknowledged:
First, given the rapid development of the field, some very recent advances may not be included. Additionally, due to space constraints, we could not exhaustively cover all existing techniques and applications.
Second, our categorization of approaches into preference memory management, personalized generation and rewarding, and alignment through feedback, while useful for organization, may oversimplify the complex interrelationships between these components.
Third, the evaluation methods and metrics discussed in this survey largely reflect current practices, which may not fully capture the nuanced aspects of personalization quality. The field lacks standardized evaluation frameworks, making it challenging to compare different approaches objectively.
Finally, while we attempted to include diverse perspectives on ethical considerations and societal impacts, our discussion may not comprehensively address all potential implications of personalized alignment, particularly in emerging application scenarios.